C. How many centimeters are in a meter?

D. What is the mode of the data set below:

5;6;1;4;7;3;8;2;4;3;8;9;4;6;1;2;6;5

E. Calculate the area of the darker shaded outer part below.The larger rectangle is 15cm long and 9 cm wide, the sides of the smallest square are 5cm each.


1
Expert's answer
2022-06-05T10:10:25-0400

C.

There are 100 centimeters in a meter.


D.

1, 1, 2, 2, 3, 3,4, 4, 4, 5, 5, 6, 6, 6, 7, 8, 8, 9

Mode are 6, 4, each appeared 3 times.


E.


"Area=15(9)\u2212(5)^2=110(cm ^2)"